• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Zolnierczuk

    Log entry time 16:09:42 on June14,2001

    Entry number 63081

    keyword=day shift summary


    8:05 attempted 5uA beam on the BeO target, beam tripping every ~40 secs
    MCC says that they are not able to deliver anything above 1.5-3uA
    After discussion with RC and spokesperson we decided that MCC should
    work on the problem - their estimate is ~1hr
    put in empty target
    9:30 EPR (38.2)
    9:50 beam back, MCC doing tests (got 15uA at some point)
    10:30 BeO target in position
    10:35 beam available beam at (-2.5,0); Raster check, looks good (4mm)
    10:50 fast feedback on, pol3He target in beam
    10:55 RUN 1549/20549 short test run
    11:10 RUN 1550/20550 production
    12:23 RUN 1551/20551 production
    12:30 NMR (34.8%)
    12:36 beam down, magnet thermal interlock problem (MCC says at least 1hr)
    13:40 still the magnet problem (MCC says 4pm the earliest)
    13:45 EPR done (30.7%)
    13:50 MCC wants to check the OTR - we are in controlled access
    14:10 Compton guy wants to go the Hall to check the electronics
    14:15 Nilanga called - the Big Bosses are concerned about the polarization loss
    we are going to the transverse mode
    14:30 Xiaochao goes in to the laser hut to do some work (change fiber,
    measure power output of laser #5, put some label)
    16:00 end of shift no beam (since 12:36)